# Franca: Nested Matryoshka Clustering for Scalable Visual Representation Learning

### v1

@@ 1 Introduction
We achieve strong results on dense prediction tasks—including in-context learning, linear segmentation, and overclustering—outperforming DINOv2-G by up to 3%. We also surpass DINOv2-G by 4% on average across five OOD detection benchmarks, improve 3D understanding, and match its performance on classification, all without relying on proprietary data.
@@ 2 Related Work
The reliance on proprietary datasets in the training of current vision foundation models raises critical concerns regarding transparency, reproducibility, and the disentanglement of contributions. Models such as SEER (Goyal et al., 2019), DINOv2 (Oquab et al., 2024), CLIP (Radford et al., 2021), and billion-scale MAE (Singh et al., 2023) are all trained on proprietary data. This practice makes it challenging for the research community to isolate the true impact of model’s novelty and training strategies from the unique characteristics and biases of the datasets themselves. The lack of access to these datasets hinders independent verification, fair comparison, and a comprehensive understanding of what truly drives model performance. Web-SSL (Fan et al., 2025) extends the study of large-scale self-supervised pretraining by training models on publicly available MetaCLIP-2B (Xu et al., 2024) dataset, showing that models trained on open data can approach the performance of those trained on proprietary data on VLM tasks.
@@ 3 Method
We propose Franca, a scalable open-source self-supervised learning framework built on iBoT (Zhou et al., 2022b) and pretrained on large public image datasets. Franca tackles key limitations in existing vision SSL models through three main components. First, we use CyclicMask inspired from (Darcet et al., 2024), a masking strategy that circularly shifts masked patches to break simple spatial continuity and promote the learning of semantic features. Second, we introduce Matryoshka embeddings (Kusupati et al., 2022), a nested multi-head clustering approach that shares projection layers to generate compressed multi-resolution representations; and finally, a lightweight post-pretraining step that identifies and removes feature components correlated with absolute patch positions resulting in spatially invariant representations. Figure 2 shows that each of these components provides consistent gains in both in-context segmentation and linear classification performance on standard benchmarks.
@@ 3.1 Matryoshka Representations for Efficient Multi-Granular Learning
Our framework supports hierarchical learning: coarse heads capture global semantics, while fine heads focus on local structure akin to early clustering works (Ji et al., 2019; Asano et al., 2020; Van Gansbeke et al., 2020) and unlike most recent representation learning works that optimize only a single feature space (Oquab et al., 2024; Tschannen et al., 2025; Radford et al., 2021). As shown in Figure 3, Franca significantly outperforms DINOv2 (Oquab et al., 2024) across all embedding sizes, especially under heavy compression. For fairness, we note that DINOv2 was not trained for dimensional truncation; its information is spread uniformly across the feature space. Even more notably, Figure 2 shows that the largest gains (4%) from our Matryoshka framework occur in dense prediction tasks.
@@ 3.2 Balancing Spatial Distribution of Visible Patches with CyclicMask
Masked image modeling (MIM) is a core component in many self-supervised vision frameworks (Oquab et al., 2024; Zhou et al., 2022b, c), where a portion of input patches are masked and the model learns to predict the unmasked regions. Commonly adapted strategies include random masking and block masking, where patches are masked randomly or as a block across the image. While simple and stochastic, this approach lacks spatial structure, often leading to fragmented visible regions that provide limited contextual coherence as shown in Figure 4 (a) and (b).
@@ 3.3 RASA: Removal of Absolute Spatial Attributes
We first study the presence of spatial biases in the projection head that clusters the patches. To this end, we run the frozen model across images from Pascal VOC and COCO and keep track of the argmax cluster predictions for each patch location. Next, we compute the entropy across patch locations for each cluster, shown in Fig. 5. Notably, we find several semantically agnostic clusters, that instead of activating for certain content, only fire at specific image locations (see Fig. 5, left). More generally, we find that clusters tend to mix semantics and spatial position (e.g., a cluster for a car tire on the right vs another cluster for the same on the left of the image), indicating that even DINOv2 contains undesirable mixing of content and geometry. In contrast, Franca and RASA significantly increase the entropy of patch locations for almost all clusters, indicating better separation of content and geometry.
To address this, we propose Removal of Absolute Spatial Attributes (RASA), a post-training strategy to disentangle position information from patch embeddings. Once our model is pretrained, we use patch features to suppress the dimensions of the embedding space that are predictive of patch location.
@@
As each iteration of RASA in Eq. 2 can be parameterised as a linear matrix-multiply with as follows,
@@ 4 Implementation details
Franca’s architecture follows DINOv2 (Oquab et al., 2024) without registers, using Vision Transformers (Dosovitskiy et al., 2021) of varying model capacities: ViT-B with 86M parameters, ViT-L with 300M, and ViT-G with 1.1B. All models are trained from scratch for 500 epochs (625K iterations) without distillation from larger models, unlike DINOv2, which distills from ViT-G into smaller variants. We employ Matryoshka (Kusupati et al., 2022) with five nested heads on the ViT backbone with feature dimension .
For LAION-600M, we use global crops scale of , following DINOv2-style augmentations. Stochastic depth regularization is set to 0.1 for ViT-B and 0.4 for ViT-L and ViT-G. We use a total batch size of 2048 for ViT-B and 3072 for both ViT-L and ViT-G, distributed across 32, 64 and 128 H100 GPUs for ViT-B, ViT-L and ViT-G respectively. The learning rate is set to for the Base model and for the Large and Giant variants, using a cosine schedule with warmup of 80 epochs.
@@ 5.1 Image Classification
We compare self-attention maps from the final layer’s [CLS] token of DINOv2, DINOv2R (with registers) (Darcet et al., 2024) and Franca in Figure 7. DINOv2 often fails to localize objects—especially under clutter or occlusion—while DINOv2R offers only minor improvements. In contrast, Franca yields sharply focused attention maps aligned with object boundaries, even for small or partially occluded instances. This suggests that our Matryoshka-style multi-head clustering promotes semantically rich features and finer-grained representations.
@@
Franca (ours) | ViT-L/14 | IN-21K | ✗ | 82.6 | 84.2 | 88.3 | 77.4 | 71.0 | 72.8 | 58.4 |
Franca (ours) | ViT-L/14 | LAION-600M | ✗ | 82.0 | 83.8 | 87.0 | 76.5 | 71.3 | 72.0 | 56.9 |
@@
Franca (ours) | ViT-G/14 | IN-21K | ✗ | 83.1 | 85.9 | 88.7 | 77.8 | 75.3 | 78.6 | 62.2 |
Franca (ours) | ViT-G/14 | LAION-600M | ✗ | 82.8 | 85.0 | 87.0 | 77.2 | 74.8 | 78.1 | 62.2 |
@@
We evaluate the generalization ability of Franca under natural distribution shifts using linear probing on ImageNet-A (Hendrycks et al., 2021a), ImageNet-R (Hendrycks et al., 2021b), and Sketch (Wang et al., 2019), which retain the original label space but introduce semantic or stylistic variations. As shown in Table 2, Franca demonstrates strong robustness. For instance, Franca-G matches the performance of DINOv2-G across all three datasets and outperforms OpenCLIP-G by 11.5% on ImageNet-A, despite OpenCLIP being trained on over 30 more data.
@@
Franca (ours) | ViT-B/14 | 91.6 | 99.0 | 92.0 | 77.0 | 88.7 | 79.3 | 82.2 | 93.8 | 96.2 | 99.7 | 88.3 | 89.8 |
Franca (ours) | ViT-L/14 | 94.3 | 99.0 | 93.7 | 78.9 | 89.5 | 81.3 | 83.8 | 96.8 | 97.4 | 99.5 | 90.9 | 91.4 |
@@ 5.3 Probing 3D awareness
As shown in Table 6, Franca achieves strong performance across both tasks, outperforming DINOv2 (reproduced on IN-21K) and exceeding Open-CLIP and SigLIPv2 baselines of similar scale. On SPair-71k, using an input resolution of , Franca outperforms DINOv2 distilled from DINOv2-G and pretrained on LVD142M by 3% on ViT-L and 1.5% on ViT-G. Notably, Franca maintains high correspondence fidelity even for large viewpoint changes (), suggesting robustness to geometric deformations.

### v2

@@ 1 Introduction
We achieve strong results on dense prediction tasks, including in-context learning, linear segmentation, and overclustering, outperforming DINOv2-G by up to 3%. We also surpass DINOv2-G by 4% on average across five OOD detection benchmarks, improve 3D understanding, and match its performance on classification, all without relying on proprietary data.
@@ 2 Related Work
The reliance on proprietary datasets in the training of current vision foundation models raises critical concerns regarding transparency, reproducibility, and the disentanglement of contributions. Models such as SEER (Goyal et al., 2019), DINOv2 (Oquab et al., 2024), CLIP (Radford et al., 2021), and billion-scale MAE (Singh et al., 2023) are all trained on proprietary data. This practice makes it challenging for the research community to isolate the true impact of model’s novelty and training strategies from the unique characteristics and biases of the datasets themselves. The lack of access to these datasets hinders independent verification, fair comparison, and a comprehensive understanding of what truly drives model performance. Inspired by the success and large-scale utilization of CLIP, OpenCLIP (Ilharco et al., 2021; Cherti et al., 2023) and MetaCLIP Xu et al. (2024) are the first to release models trained on public data. OpenCLIP (Cherti et al., 2023; Nezhurina et al., 2025) stands out with the fully reproducible pipeline leveraging the ready-to-use LAION dataset (Schuhmann et al., 2022; LAION, 2024), study of scaling laws and release of intermediate checkpoints. On vision-only foundation models, the performance of DINOv2 has proven difficult to match by public data models. Web-SSL (Fan et al., 2025) extends the study of large-scale self-supervised pretraining by training models on publicly available MetaCLIP-2B (Xu et al., 2024) dataset, showing that models trained on open data can approach the performance of those trained on proprietary data on VLM tasks, but still below DINOv2 on visual perception tasks.
@@ 3 Method
We propose Franca, a scalable open-source self-supervised learning framework built on iBoT (Zhou et al., 2022a) and pretrained on large public image datasets. Franca tackles key limitations in existing vision SSL models through three main components. First, we use CyclicMask inspired from (Darcet et al., 2024), a masking strategy that circularly shifts masked patches to break simple spatial continuity and promote the learning of semantic features. Second, we introduce Matryoshka embeddings (Kusupati et al., 2022), a nested multi-head clustering approach that shares projection layers to generate compressed multi-resolution representations; and finally, a lightweight post-pretraining step that identifies and removes feature components correlated with absolute patch positions resulting in spatially invariant representations. Figure 2 shows that each of these components provides consistent gains in both in-context segmentation and linear classification performance on standard benchmarks.
@@ 3.1 Matryoshka Representations for Efficient Multi-Granular Learning
Our framework supports hierarchical learning: coarse heads capture global semantics, while fine heads focus on local structure akin to early clustering works (Ji et al., 2019; Asano et al., 2020; Van Gansbeke et al., 2020) and unlike most recent representation learning works that optimize only a single feature space (Oquab et al., 2024; Tschannen et al., 2025; Radford et al., 2021). As shown in Figure 4, Franca significantly outperforms DINOv2 (Oquab et al., 2024) across all embedding sizes, especially under heavy compression. For fairness, we note that DINOv2 was not trained for dimensional truncation; its information is spread uniformly across the feature space. Even more notably, Figure 2 shows that the largest gains (4%) from our Matryoshka framework occur in dense prediction tasks. This is further supported by the PCA visualization in Figure 3 where Franca retains coherent part-leve structure well beyond its trained dimensions, unlike DINOv2, which looses semantic alignment.
@@ 3.2 Balancing Spatial Distribution of Visible Patches with CyclicMask
Masked image modeling (MIM) is a core component in many self-supervised vision frameworks (Oquab et al., 2024; Zhou et al., 2022a, b), where a portion of input patches are masked and the model learns to predict the unmasked regions. Commonly adapted strategies include random masking and block masking, where patches are masked randomly or as a block across the image. While simple and stochastic, this approach lacks spatial structure, often leading to fragmented visible regions that provide limited contextual coherence as shown in Figure 5 (a) and (b).
@@ 3.3 RASA: Removal of Absolute Spatial Attributes
We first study the presence of spatial biases in the projection head that clusters the patches. To this end, we run the frozen model across images from Pascal VOC and COCO and keep track of the argmax cluster predictions for each patch location. Next, we compute the entropy across patch locations for each cluster, shown in Figure 6. Notably, we find several semantically agnostic clusters, that instead of activating for certain content, only fire at specific image locations (see Figure 6, left). More generally, we find that clusters tend to mix semantics and spatial position (e.g., a cluster for a car tire on the right vs another cluster for the same on the left of the image), indicating that even DINOv2 contains undesirable mixing of content and geometry. In contrast, Franca and RASA significantly increase the entropy of patch locations for almost all clusters, indicating better separation of content and geometry.
To address this, we propose Removal of Absolute Spatial Attributes (RASA), a post-training strategy to disentangle position information from patch embeddings. Once our model is pretrained, we use patch features to suppress the dimensions of the embedding space that are predictive of patch location. For brevity, we will use to refer to throughout this section.
@@
As each iteration of RASA in Equation 2 can be parameterised as a linear matrix-multiply with as follows,
@@ 4 Implementation details
Franca’s architecture follows DINOv2 (Oquab et al., 2024) without registers, using Vision Transformers (Dosovitskiy et al., 2021) of varying model capacities: ViT-B with 86M parameters, ViT-L with 300M, and ViT-G with 1.1B. All models are trained from scratch for 625K iterations without distillation from larger models, unlike DINOv2, which distills from ViT-G into smaller variants. We employ Matryoshka (Kusupati et al., 2022) with five nested heads on the ViT backbone with feature dimension .
For LAION-600M, we use global crops scale of , following DINOv2-style augmentations. Stochastic depth regularization is set to 0.1 for ViT-B and 0.4 for ViT-L and ViT-G. We use a total batch size of 2048 for ViT-B and 3072 for both ViT-L and ViT-G, distributed across 32, 64 and 128 H100 GPUs for ViT-B, ViT-L and ViT-G respectively. The learning rate is set to for the Base model and for the Large and Giant variants, using a cosine schedule with warmup of 100K iterations.
@@ 5.1 Image Classification
We compare self-attention maps from the final layer’s [CLS] token of DINOv2, DINOv2R (with registers) (Darcet et al., 2024) and Franca in Figure 8. DINOv2 often fails to localize objects, especially under clutter or occlusion, while DINOv2R offers only minor improvements. In contrast, Franca yields sharply focused attention maps aligned with object boundaries, even for small or partially occluded instances. This suggests that our Matryoshka-style multi-head clustering promotes semantically rich features and finer-grained representations.
@@
Franca (ours) | ViT-L/14 | IN-21K | ✗ | 82.2 | 84.5 | 88.3 | 77.4 | 71.0 | 72.8 | 58.4 |
Franca (ours) | ViT-L/14 | LAION-600M | ✗ | 81.9 | 84.4 | 87.0 | 76.5 | 71.3 | 72.0 | 56.9 |
@@
Franca (ours) | ViT-G/14 | IN-21K | ✗ | 83.0 | 85.9 | 88.7 | 77.8 | 75.3 | 78.6 | 62.2 |
Franca (ours) | ViT-G/14 | LAION-600M | ✗ | 81.2 | 85.0 | 87.0 | 77.2 | 74.8 | 78.1 | 62.2 |
@@
We evaluate the generalization ability of Franca under natural distribution shifts using linear probing on ImageNet-A (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), and Sketch (Wang et al., 2019), which retain the original label space but introduce semantic or stylistic variations. As shown in Table 2, Franca demonstrates strong robustness. For instance, Franca-G matches the performance of DINOv2-G across all three datasets and outperforms OpenCLIP-G by 11.5% on ImageNet-A, despite OpenCLIP being trained on over 30 more data.
@@
Franca (ours) | ViT-B/14 | 90.6 | 98.7 | 90.9 | 77.0 | 88.7 | 75.2 | 81.7 | 94.1 | 96.2 | 99.7 | 86.2 | 88.9 |
Franca (ours) | ViT-L/14 | 94.3 | 99.4 | 94.1 | 79.9 | 89.5 | 81.3 | 84.1 | 95.1 | 97.4 | 99.8 | 91.1 | 91.5 |
@@ 5.3 Probing 3D awareness
As shown in Table 6, Franca demonstrates strong performance across both tasks, outperforming DINOv2 (reproduced on IN-21K) and surpassing Open-CLIP and SigLIPv2 baselines of similar scale. On SPair-71k, using an input resolution of , Franca outperforms DINOv2 distilled from DINOv2-G and pretrained on LVD142M by 3% on ViT-L and 1.5% on ViT-G. On a higher-resolution of , Franca outperforms DINOv2 at different scales when using the same pretraining dataset (Imagenet-21K) without any distillation. Notably, Franca maintains high correspondence fidelity even for large viewpoint changes (), suggesting robustness to geometric deformations.